OpenStudy (anonymous):

how do i simplify 4x^2-7x+3 divided by x^2+5x-6?

OpenStudy (ash2326):

we have \[\frac{4x^2-7x+3}{x^2-5x+6}\] let's divide (x^2-5x+6)*4=4x^2-20x+24 rewriting the numerator as 4x^2-7x+6=(4x^2-20x+24)+13x-18 so we have now \[\frac{(4x^2-20x+24)+13x-18 }{x^2-5x+6}\] if we simplify we get \[4+\frac{13x-18 }{x^2-5x+6}\]

OpenStudy (anonymous):

is there more simplifying? because thats not one of my options

OpenStudy (anonymous):

i wrote the question wrong, the denomenator is supposed to be x^2+5x-6

OpenStudy (ash2326):

then 4(x^2+5x-6)=4x^2+20x-24 so 4x^2-7x+3=(4x^2+20x-24)-27x+27 so we get \[4+\frac{-27x+27}{x^2+5x-6}\]

OpenStudy (anonymous):

these are my options

OpenStudy (anonymous):

\[x^2 + 5x - 6 = (x-1)(x+6)\] \[4x^2-7x+3 = 4x^2-4x-3x+3=4x(x-1)-3(x-1)=(4x-3)(x-1)\] So cancelling (x-1) you can see the answer

OpenStudy (anonymous):

which answer is it?

OpenStudy (anonymous):

\[\frac{4x-3}{x+6}\]